# A Lightweight and Gradient-Stable Neural Layer

### v3

@@ Abstract
We propose a neural-layer architecture based on Householder weighting and absolute-value activating, hence called Householder-absolute neural layer or simply Han-layer. Compared to a fully connected layer with $d$-neurons and $d$ outputs, a Han-layer reduces the number of parameters and the corresponding complexity from $O(d^{2})$ to $O(d)$. The Han-layer structure guarantees two desirable properties: (1) gradient stability (free of vanishing or exploding gradient), and (2) 1-Lipschitz continuity. Extensive numerical experiments show that one can strategically use Han-layers to replace fully connected (FC) layers, reducing the number of model parameters while maintaining or even improving the generalization performance. We will also showcase the capabilities of the Han-layer architecture on a few small stylized models, and discuss its current limitations.
@@ 1 Introduction
The advancement of neural networks has sparked a revolution across multiple disciplines. However, with the increasing complexity of the problems they are designed to solve, the models have become more cumbersome and computationally expensive. For example, the cost of training a large language model (LLM) can be in millions energyconsider , and the number of parameters of such a model is often several billion. On the other hand, small-scale devices, such as mobile or vehicle-mounted devices, are resource-constrained, making it hard to deploy such colossal models. Consequently, there is a growing demand for leaner models that have fewer parameters and can operate more efficiently energyconsider ; vitsurvey . Researchers have employed methods such as devising low-complexity operators, and compressing and pruning models, as demonstrated in recent surveys vitsurvey ; liang2021pruning .
So far, however, there is still a lack of widely used low-complexity layer architecture that can effectively replace fully connected layers (under suitable conditions, of course). In this paper, we propose a new neural network layer called the Householder absolute-value neural layer (or Han-layer), which replaces the dense weight matrix in a fully connected layer with a Householder matrix and uses the absolute-value function as the activation function. Our choice of Householder matrix is motivated by two key factors. First, it is an orthogonal matrix, ensuring the stability of gradients during training. Second, a $d\times d$ Householder matrix involves just $d$ parameters, thereby making it possible to yield a substantial reduction in the overall parameter count.
Furthermore, our selection of ABS as the activation function is guided by the following considerations. Firstly, the Jacobian matrix of ABS, when applied element-wise, is an orthogonal matrix with ±$\pm 1$ on its diagonal. By construction, deep network structures built with multiple Han-layers demand relatively few parameters and theoretically avoid the gradient vanishing or exploding problem. Therefore, our model obviates the need for employing normalization techniques and conventional residual connections batchnorm ; weightnorm ; resnet ; residualnormpre . Additionally, ABS, as a piece-wise linear function, which can be viewed as a combination of two ReLU functions, has very low computational complexity.
@@ 1.1 Contributions
We propose a new, lightweight and gradient stable layer architecture called Han-layer, and carefully examine its properties. We conduct extensive experiments to evaluate the capabilities of neural networks equipped with multiple Han-layers (or simply HanNets for brevity). In addition, since Han-layers are 1-Lipschitz continuous, they are naturally resistant to adversarial attacks to a degree.
@@
    On several standard datasets for regression and image classification, our experiments indicate that HanNets can significantly reduce the number of model parameters, even within models already characterized by lightweight, while maintaining, and in some cases improving, generalization performance.
@@
    On stylized small problems (see Section 4), our experiments show the existence of some structured data on which HanNets can greatly outperform conventional Multi-layer Perceptrons (MLPs) in terms of generalization, as is demonstrated in Figure 1 where a HanNet shows remarkable superiority over traditional FCNets and produces a nearly perfect result.
@@
Figure 1: Landscapes and top views from FC and Han models on the checkerboard set. (a) for FCNet, (b) for HanNet.
Even though our empirical results in this paper are not obtained from huge-scale models, they should be sufficient to establish that the new Han-layer architecture represents a useful technique to be included in the existing toolbox of deep learning.
@@ 1.2 Notations
We define a function $F_{L}$ from $\mathbb{R}^{m}\to\mathbb{R}^{n}$ realized by a deep neural network with depth $L$,
@@
where $\textbf{W}=\{W_{1},...,W_{L}\}$ and $\textbf{b}=\{b_{1},...,b_{L}\}$ is the collections of weight matrices and bias vectors, $\psi_{i}$ is the layer function, and if $\psi_{i}$ is a fully connected layer with activation function, then
@@
where $W_{i}$ is the weight matrix and $\phi(\cdot)$ is the element-wise activation function. The deep neural network (DNN) function $F_{L}(x,\textbf{W},\textbf{b})$ can be computed via forward propagation,
@@
For the convenience of subsequent analysis, here we calculate the typical Jacobian matrix involved in the gradient computation through backpropagation, denoted as the $G$-matrix,
@@
where $\nabla\phi(z_{k})$ is the diagonal, Jacobian matrix of $\phi(\cdot)$ evaluated component-wise at a vector $z_{k}$. The spectrum of a so-called $G_{L}$-matrix characterizes whether vanishing or exploding gradient would happen or not. Roughly speaking, $G_{L}$-matrix is the dominant part of the gradient. As $L\rightarrow\infty$, vanishing gradient corresponds to $\|G_{L}\|\rightarrow 0$ and exploding gradient to $\|G_{L}\|\rightarrow\infty$.
@@ 1.3.1 Absolute-Value Function for Activation
In earlier periods of neural network research, piece-wise linear model structures were considered that involved ABS functions batruni1991multilayer ; lin1992canonical . However, ABS failed to gain mainstream popularity and was seldom utilized, e.g. not mentioned in this survey jagtap2022important . Instead, ReLU has become the most widely used activation function in recent years, as demonstrated in nair2010rectified . The author in karnewar2018aann suggests using ABS activation in a “bidirectional neuron” architecture based on interpretability considerations. More recently, in zhang2021variability , the authors show that under the same conditions, ABS can better resist the occurrence of “collapse to constant” and maintain network variability better than ReLU. Additionally, in beknazaryan2021analytic , the author demonstrates that neural networks with absolute value activation functions can $\epsilon$-approximate functions that are analytic on certain regions. For more information about piecewise linear activations including the absolute-value function, see this survey tao2022piecewise .
1.3.2 Orthogonal Weight Matrix
Orthogonal weight-matrix initialization has been the subject of theoretical and empirical investigations and has proven to be useful in deep learning, that is, orthogonal initial weights can speed up convergence relative to standard Gaussian initial weights, as evidenced by studies such as Hu2020Provable ; huang2018orthogonal ; le2015simple ; pennington2018emergence ; xiao2018dynamical . Additionally, this orthogonal weight-matrix approach has been applied to recurrent and convolution neural networks to improve model performance, as shown in studies such as householderrnnefficient ; wang2020orthogonal ; arjovsky2016unitary ; jia2019orthogonal ; tomczak2016improving ; vorontsov2017orthogonality ; wisdom2016full ; zhang2018stabilizing ; householderactivation . For example, the authors in jia2019orthogonal show that orthogonal neural networks have better generalization performance based on the local isometry property.
In this article, we mainly take advantage of orthogonality for gradient stability. In a neural network mapping from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, if the Jacobian matrix of each layer function is orthogonal, then the total Jacobian matrix of the mapping will be orthogonal since products of orthogonal matrices remain orthogonal. Therefore, with all eigenvalues maintaining unit modulus, there will be no gradient explosion or vanishing.
Moreover, a simple method brock2017neural to encourage near-orthogonal weight matrices in FCNets is to regularize each weight matrix $W$ by the term:
$\|W^{{\intercal}}W-I\|_{F}^{2}\ \ \mathrm{or}\ \ \|WW^{{\intercal}}-I\|_{F}^{2},$
where $I$ denotes the identity matrix. For a comparison purpose, we will use this regularization in some subsequent experiments, see Table 3.
1.3.3 Householder Weighting
The Householder reflection matrix associated with a nonzero vector $u\in\mathbb{R}^{d}$ is
$H(u)=I-2\frac{uu^{{\intercal}}}{u^{{\intercal}}u},$ (5)
where $I$ denotes the identity matrix in $\mathbb{R}^{d}$. This matrix is both symmetric and orthogonal. In addition, the number of parameters in $H(u)$ is an order of magnitude smaller than that of a general dense matrix.
In some works such as the orthogonal RNN (oRNN) householderrnnefficient , products of Householder reflection matrices are used to represent transition matrices to alleviate issues related to vanishing and exploding gradients. A more recent work in this direction is mathiasen2020one .
@@ 2 Introduction to Han-layer
Figure 2: Visualization of An MLP-layer with ReLU activation function and a Han-layer. The $W$ is a weight matrix, $u$ is a nonzero vector, $\bigoplus$ denotes element-wise addition and ABS is the absolute-value function.
@@ 2.1 Absolute-value Activating
Nonlinearity, continuity and differentiability (almost everywhere) are universally required for activation functions with obvious or well-explained rationales groupsort . Hence, conditions A1-A2 are natural.
@@
Let ϕ:normal-:italic-$\phi:\mathbb{R}\rightarrow\mathbb{R}$ have Property A and $|\mathcal{C}_{\phi}|$ denote the cardinality of $\mathcal{C}_{\phi}$. Then $|\mathcal{C}_{\phi}|\geq 1$.
@@
Let ϕ:normal-:italic-$\phi:\mathbb{R}\rightarrow\mathbb{R}$ have Property A. Then $\phi$ is piecewise linear so that each piece has the functional form $\phi(t)=\pm t+\mathrm{const}$ with a different constant . In the case of the minimum cardinality $|\mathcal{C}_{\phi}|=1$ with $\mathcal{C}_{\phi}=\{\alpha\}$, $\phi$ has the functional form
@@ 4 Stylized Datasets
Figure 5: Checkerboard datasets. The right figure is the training set (25%).
@@
(a) FCNet: the best test accuracy is up to 87%.
@@ 4.1 A preliminary explanation
To explain HanNet’s exceptional performance on the checkerboard dataset, we take an intuitive viewpoint based on the amount of nonlinearity provided by different models. In most feed-forward networks, the main source of nonlinearity is from the applications of nonlinear activation functions. Given an MLP-layer that maps from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, the number of activation-function applications is $d$, with $d^{2}$ parameters in weight matrix $W$. Consequently, the activation-function utilization rate is $1/d$ (i.e., $d$ versus $d^{2}$), which serves as a normalized measure of nonlinearity. In contrast, with $d$ parameters in $u$, our Han-layer has an activation-function utilization rate 1 (i.e., $d$ versus $d$). Therefore, several Han-layers together can potentially produce more nonlinearity than a single MLP layer can, while using far fewer parameters.
It is reasonable to argue that with more nonlinearity a HanNet would be better equipped than an FCNet to approximate complex shapes as in the checkerboard dataset and capture the underlying patterns. On the other hand, an FCNet, characterized by over-parameterization, may be able to readily fit training data by memorizing them, but still fail to fully discern the underlying patterns. This seems to explain well the observed generalization discrepancy in Figure 1.
@@ 4.2 Ablation Study on Checkerboard
We conducted an ablation study on a $100\times 20$ network framework with seven different configurations to investigate how much Householder weighting and ABS activating contribute to the remarkable results on the checkerboard dataset. Table 3 lists the results, where we used either Householder (HH) or fully connected (FC) layers with/without orthogonal regularization, and either ABS, ReLU, MaxMin or GroupSort groupsort activation functions. The latter two can keep the gradient norm unchanged groupsort . We obtained orthogonal FCNets (OrthFC) by regularizing each weight matrix $W$ as brock2017neural :
@@ 4.3 Another Stylized Dataset
The remarkable performance of HanNet on the checkerboard dataset is not an isolated case, as shown by the experiment in hong2022activation . Here, we fit another function $g(x)$ using HanNet and FCNet, defined by the sum of two sine-products (of low and high frequencies)
@@
Figure 7: Landscapes and top views from FC and Han models. From left to right: the original data, HanNet, FC-ReLU, FC-ABS. The test loss by HanNet is 0.034 while it is about 0.37 by the other two models.

### v4

@@ Abstract
To enhance resource efficiency and model deployability of neural networks, we propose a neural-layer architecture based on Householder weighting and absolute-value activating, called Householder-absolute neural layer or simply Han-layer. Compared to a fully connected layer with $d$-neurons and $d$ outputs, a Han-layer reduces the number of parameters and the corresponding computational complexity from $O(d^{2})$ to $O(d)$. The Han-layer structure guarantees that the Jacobian of the layer function is always orthogonal, thus ensuring gradient stability (i.e., free of gradient vanishing or exploding issues) for any Han-layer sub-networks. Extensive numerical experiments show that one can strategically use Han-layers to replace fully connected (FC) layers, reducing the number of model parameters while maintaining or even improving the generalization performance. We will also showcase the capabilities of the Han-layer architecture on a few small stylized models, and discuss its current limitations.
@@ 1 Introduction
The advancement of neural networks has sparked a revolution across multiple disciplines. However, as models grow larger and larger, so do their demands for resources such as energy and computing costs. For example, training a large language model (LLM), with the number of parameters reaching hundreds of billions, can cost millions of U.S. dollars [1]. On the other hand, most devices in real applications are resource-constrained, such as mobile or vehicle-mounted devices, making it hard to deploy over-sized models. Consequently, there is a growing demand for leaner models with fewer parameters that can operate efficiently with adequate functionality [1, 2].
Although researchers have investigated various methods, such as model compressing and pruning, to reduce model sizes (see recent surveys [2, 3] and references therein), so far there is still a lack of a low-complexity (in term of parameter and operation counts) and gradient-stable layer architecture that can effectively replace fully connected layers under suitable conditions. In this paper, we propose a new neural network layer called the Householder Absolute-value Neural layer (or Han-layer), which replaces the dense weight matrix in a fully connected layer with a Householder matrix and uses the absolute-value function as the activation function.
Our choice of Householder matrix is motivated by two key factors. First, it is an orthogonal matrix, ensuring gradient stability during training. Second, a $d\times d$ Householder matrix involves just $d$ parameters, thereby making it possible to yield a substantial reduction in the overall parameter count even when multiple Han-layers are used to replace a single fully connected layer. On the other hand, our selection of absolute-value function, or ABS, as the activation function is guided primarily by the consideration that the Jacobian matrix of ABS, applied element-wise, is a diagonal and orthogonal matrix (with ±$\pm 1$ on its diagonal). Additionally, ABS is a piecewise linear function, like the popular ReLU function, with a low computational complexity.
It is crucial to underscore the complementary roles played by Householder weighting and ABS activating. The combination of the two ensures orthogonality of the Jacobian matrices for each and all layer functions (see Equation (4) below), regardless of the number of layers. This orthogonality ensures gradient stability without needing to deploy conventional techniques such as normalizations and residual connections [4, 5, 6, 7]. In addition, this very combination inherits the benefits of being lightweight (from Householder weighting) and being computationally efficient (from ABS activating).
@@ 1.1 Contributions
We propose a new, lightweight and gradient-stable layer architecture called Han-layer, and carefully examine its properties. We conduct extensive experiments to evaluate the capabilities of neural networks equipped with multiple Han-layers (or simply HanNets for brevity). In addition, since Han-layers are 1-Lipschitz continuous, they are naturally resistant to adversarial attacks to a degree.
@@
    On several standard datasets for regression and image classification, our experiments indicate that HanNets can significantly reduce the number of model parameters, even for models already characterized as lightweight, while maintaining, and in some cases improving, generalization performance.
@@
    On stylized small problems (see Section 4), our experiments show the existence of some structured data on which HanNets can greatly outperform conventional Multi-layer Perceptrons (MLPs) in terms of generalization, as is demonstrated in Figure 1 where the HanNet result is nearly perfect.
Refer to captionRefer to caption
@@
Figure 1: Landscapes and top views for FC and Han models on the checkerboard dataset: (a) FCNet, and (b) HanNet.
Even though our empirical results in this paper [^1] For reproducibility, all the relevant codes are posted in GitHub at the link: https://github.com/yyy32/HanNet. are obtained from small- to moderate-scale models, they should be sufficient to demonstrate that the new Han-layer architecture represents a useful technique to be included in the existing toolbox of deep learning. It may be particularly useful in building lightweight models, for example, in mobile applications.
@@ 1.2 Notations
We define a function $F_{L}$ from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$ realized by a deep neural network with depth $L$ as follows
@@
where $\textbf{W}=\{W_{1},...,W_{L}\}$ and $\textbf{b}=\{b_{1},...,b_{L}\}$ are the collections of weight matrices and bias vectors, respectively; $\psi_{i}$ is the $i$-th layer function with an activation function $\phi$, defined as
@@
where the scalar function $\phi(\cdot)$ is applied element-wise. The deep neural network (DNN) function $F_{L}(x,\textbf{W},\textbf{b})$ can be computed via forward propagation,
@@
For convenience of subsequent analysis, here we calculate the following Jacobian matrix involved in the gradient computation through back-propagation, denoted as the $G_{L}$-matrix,
@@
where $W_{k}\nabla\phi(z_{k})$ is the Jacobian matrix of the $k$-th layer, and $\nabla\phi(z_{k})$ is the diagonal, Jacobian matrix of $\phi(\cdot)$ evaluated component-wise at a vector $z_{k}$. The spectral norm of the $G_{L}$-matrix can serve as an indication whether gradient vanishing or exploding will happen. Asymptotically speaking, as $L\rightarrow\infty$, gradient vanishing happens if $\|G_{L}\|_{2}\rightarrow 0$ and gradient exploding happens when $\|G_{L}\|_{2}\rightarrow\infty$.
We denote the Householder (reflection) matrix associated with a nonzero vector $u\in\mathbb{R}^{d}$ as
$H(u)=I-2\frac{uu^{{\intercal}}}{u^{{\intercal}}u},$ (5)
where $I$ is the identity matrix of dimension $d$. $H(u)$ is both symmetric and orthogonal. The number of parameters in $H(u)$ is $d$, an order of magnitude lower than that of a general $d$ by $d$ matrix.
@@ 1.3 Related Work
The proposed Han-layer architecture combines ABS activating with Householder weighting based on two central considerations: (1) guaranteed orthogonality of every layer Jacobian $H(u_{k})\nabla\phi(z_{k})$, and (2) low computational complexity. Consequently, the total Jacobian matrix of multiple Han-layers remains orthogonal since products of orthogonal matrices are orthogonal, ensuring gradient stability (i.e., no gradient exploding or vanishing) in any HanNet portion of a network.
Since the proposed combination of the two key components has not yet been previously investigated, in the following two subsections we will survey related work for ABS activating and Householder weighting separately. We note that neither of the two techniques has been widely used, though neither is really new.
@@ 1.3.1 Absolute-Value Function for Activation
ABS stands out as one of the simplest nonlinear functions whose Jacobian matrices, whenever existing, are always orthogonal (see Sec. 2.1). In addition, it is piecewise linear and easy to apply at a low cost. In the literature, we found only a couple of other piecewise linear activation functions with orthogonal Jacobian matrices (when existing) which are named MaxMin and GroupSort and proposed in [8]. Later on, we will compare ABS with these two in an ablation experiment, see Table 3.
@@
In earlier periods of neural network research, piecewise linear function ABS was used [9, 10]. However, ABS failed to gain mainstream popularity and was seldom utilized. For example it is not mentioned in this recent survey [11] of activation functions. Instead, ReLU has become the most widely used (piecewise linear) activation function, as demonstrated in [12]. The author in [13] suggests using ABS activation in a “bidirectional neuron” architecture based on interpretability considerations. More recently, in [14], the authors show that under the same conditions, ABS can better resist the occurrence of “collapse to constant” and maintain network variability better than ReLU. Additionally, in [15], the author demonstrates that neural networks with absolute value activation functions can $\epsilon$-approximate functions that are analytic on certain regions.
Beside ABS and ReLU, there are many other piecewise linear activation functions studied in the literature, see the survey [16] and a specific example in [17]. Most previous studies on piecewise linear activation functions are not directly related to ABS activating since Jacobian orthogonality is out of their scope of considerations.
1.3.2 Householder Weight Matrices
In recurrent neural networks (RNN), products of Householder reflection matrices are utilized to represent transition matrices in order to alleviate issues related to gradient instability, including [18, 19] (see [20] for a related variant). There are two major differences between the previous use of Householder weighting and ours: (1) we couple each Householder weight matrix with the ABS activation to ensure orthogonality of the layer Jacobian; (2) we only make use of a single Householder matrix per layer to create light-weight layers.
Beside mitigating gradient instability, the sustained orthogonality (as opposed to at initialization only) of Householder matrices has been used to enhance learning quality in a number of application scenarios, such as in generative imaging [21], variational auto-encoders [22], model robustness [23] and speaker anonymization [24].
In addition to using Householder reflection matrices, other methods are proposed to enforce sustained orthogonality, exemplified by [25, 26, 27] where Householder, permutation, Cayley Transformation and/or other forms of orthogonal or unitary matrices are utilized (or combined) to form weight matrices. Other works, such as [28, 29, 30], try to keep weight matrices approximately orthogonal by various means, including using singular-value decomposition or regularization terms such as $\|W^{{\intercal}}W-I\|_{F}^{2}$. In [29], the authors demonstrate that orthogonal neural networks, leveraging local isometry properties, manifest superior generalization performance.
Finally, orthogonal weight-matrix initializations have been the subject of theoretical and empirical investigations, such as in [31, 32, 33, 34]. Even though initialization techniques do not maintain sustained weight orthogonality during training, they have proven to be useful in deep learning.
@@ 2 Introduction to Han-layer
Figure 2: Visualization of an MLP-layer with ReLU activation and a Han-layer with ABS activation. On the left, $W$ is $d\times d$ and $Wx$ requires $O(d^{2})$ operations, while on the right $u$ is a nonzero $d$-vector and the multiplication with $x$ requires only $O(d)$ operations.
@@ 2.1 Absolute-value Activating
Nonlinearity, continuity and differentiability (almost everywhere) are universally required for activation functions with obvious or well-explained rationales [8]. Hence, conditions A1-A2 are natural.
@@
Let $\phi:\mathbb{R}\rightarrow\mathbb{R}$ have Property A and $|\mathcal{C}_{\phi}|$ denote the cardinality of $\mathcal{C}_{\phi}$. Then $|\mathcal{C}_{\phi}|\geq 1$.
@@
Let $\phi:\mathbb{R}\rightarrow\mathbb{R}$ have Property A. Then $\phi$ is piecewise linear so that each piece has the functional form $\phi(t)=\pm t+\mathrm{const}$ with a different constant . In the case of the minimum cardinality $|\mathcal{C}_{\phi}|=1$ with $\mathcal{C}_{\phi}=\{\alpha\}$, $\phi$ has the functional form
@@ 4 Stylized Datasets
Refer to captionRefer to caption
Figure 5: Checkerboard datasets. In the right figure, the dots represent the training set (25%).
@@
(a) FCNet: the best test accuracy is about 87%.
@@ 4.1 A preliminary explanation
To partially explain HanNet’s exceptional performance on the checkerboard dataset, we give an intuitive viewpoint based on the amount of nonlinearity provided by different models. In most feed-forward networks, the main source of nonlinearity is from the applications of nonlinear activation functions. Given an MLP-layer that maps from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, the number of activation-function applications is $d$, with $d^{2}$ parameters in the associated weight matrix $W$. Consequently, the activation-function utilization rate is $1/d$ (i.e., $d$ versus $d^{2}$), which serves as a measure of nonlinearity normalized with respect to the parameter count. In contrast, with $d$ parameters in $u$, our Han-layer has an activation-function utilization rate 1 (i.e., $d$ versus $d$). Therefore, several Han-layers together can potentially produce more nonlinearity than a single MLP layer can, while still using far fewer parameters. Moreover, as we discussed in Section 3.1, the absolute-value activation, coupled with Householder weighting, gives gradient stability.
It is reasonable to argue that, with high nonlinearity and gradient stability, a HanNet would be better equipped than an FCNet to approximate complex shapes as in the checkerboard dataset and capture the underlying patterns. On the other hand, an FCNet, characterized by over-parameterization, may be able to readily fit training data by memorizing them, but still fail to fully discern the underlying patterns. This seems to explain, at least partially, the observed generalization discrepancy in Figure 1.
@@ 4.2 Ablation Study on Checkerboard
We conducted an ablation study on a $100\times 20$ network framework with seven different configurations to investigate how much Householder weighting and ABS activating contribute to the remarkable results on the checkerboard dataset. Table 3 lists the results, where we used either Householder (HH) or fully connected (FC) layers with/without orthogonal regularization, and either ABS, ReLU, MaxMin or GroupSort [8] activation functions. The latter two can keep the gradient norm unchanged [8]. We obtained orthogonal FCNets (OrthFC) by regularizing each weight matrix $W$ as [30]:
@@ 4.3 Another Stylized Dataset
The remarkable performance of HanNet on the checkerboard dataset is not an isolated case, as shown by the experiment in [35]. Here, we fit another function $g(x)$ using HanNet and FCNet, defined by the sum of two sine-products (of low and high frequencies)
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: Landscapes and top views for FC and Han models. From left to right: the original data, HanNet, FC-ReLU, FC-ABS. The test loss by HanNet is 0.034 while it is about 0.37 by the other two models.